Give an example of a function which represents all types of a function. Find the composite

function (f o g) (x) given that

f = {(1,6), (4,7), (5,0)} and g = {(6,1), (7,4), (0,5)}


1
Expert's answer
2021-06-07T04:54:37-0400

Let us give an example of a function which represents all types of a function. Consider the function "h=\{(a,0),(b,1)\}." Since for "a\ne b" we have that "f(a)=0\ne 1=f(b)," we conclude that this function is injection. Taking into account that "f^{-1}(0)=\{a\}\ne \emptyset, \ f^{-1}(1)=\{b\}\ne \emptyset," we conclude that the function "f" is surjection, and hence this function is bijection.


Let us find the composite function "(f\circ g) (x)" given that "f = \{(1,6), (4,7), (5,0)\}" and "g = \{(6,1), (7,4), (0,5)\}":


"(f\circ g) (0)=f(g(0))=f(5)=0,\ (f\circ g) (6)=f(g(6))=f(1)=6,\"

"(f\circ g) (7)=f(g(7))=f(4)=7."


Therefore, "f\circ g = \{(0,0), (6,6), (7,7)\}."